# Euclid: A blue galaxy population and a brightest cluster galaxy in the making in a MaDCoWS2 galaxy cluster candidate††thanks: This paper is published on behalf of the Euclid Consortium

### 07 Mar 2026 v2

@@ 3.3.2 IRAC photometry and morphology modelling
IRAC mosaics have a pixel size of , which is six times larger than the Euclid VIS and NISP images. To measure the flux of the marginal detections in the IRAC and channels, we resample our convolved segmentation map to the correct pixel size (setting the pixels with fractional values to unity) and perform the flux and uncertainty measurements as described above.
However, at and the BCG complex is heavily blended with neighbouring galaxies. We used Galfit modelling (Peng et al., 2002, 2010a) to disentangle the structure from neighbouring galaxy fluxes. Using the high resolution of the NISP data (Fazio et al., 2004; Euclid Collaboration: Jahnke et al., 2025), we construct initial models in band. We first model the BCG complex with seven Sérsic profiles for the structure nuclei, and two additional Sérsic profiles for the diffuse emission. Because no PSF estimation is provided with Q1 frames, we approximate Euclid PSFs using a neighbouring unsaturated star. The result is shown in Fig. 6. We then make a separate model of the contaminants before merging the two models together.
Our IRAC and models are directly based on the combined -band models, with a few changes. Unlike the Euclid data, IRAC mosaics have uncertainty maps, which we added into the Galfit settings. We also delete one of the two diffuse Sérsic profiles, as we could not make the merged models converge otherwise. Likewise, we remove some of the faintest interlopers, as they were too faint to be detected in the IRAC images. For our first pass models, we only let the magnitudes change, but for our final models we allow positional changes as well to account for shifts induced by PSF miscentring; the effective radii, Sérsic indexes, axis ratios, and position angles are fixed to initial values.
@@
We measured the flux of the BCG complex in the IRAC and channels directly from the best Galfit models. To ensure consistency with the aperture photometry (see Sect. 3.3.1), we compute an aperture correction by dividing the aperture flux of the band by the Galfit flux in the same band. We find an aperture correction of .
@@ 3.3.3 Spectral energy distribution fitting
Table 2 presents the derived DECam, Euclid, and IRAC photometry, which we use to perform spectral energy distribution (SED) modelling of the entire BCG complex. We do not attempt to model the SED of individual cores: the signal-to-noise of the and DECam fluxes are not sufficient to perform a Galfit decomposition.
We adopt most of the Bagpipes (Carnall et al., 2018) standard basic assumptions (models from Bruzual and Charlot, 2003, with a Kroupa and Boily 2002 initial mass function, uniform priors), adding a dust component following the Calzetti et al. (2000) attenuation law. We replace the Bagpipes default cosmology with the Planck Collaboration: Aghanim et al. (2020) cosmological parameters. We adopt a parametric approach, testing a delayed -model and a model in which all the stars form at the same time (i.e. a burst). The burst yields a best fit SED with a smaller reduced than the delayed -model (3.34 compared to 3.75). Furthermore the characteristic time of the delayed -model is extremely short, Myr. The free parameters, priors, and results of the bursty model are presented in Table 3. The results are expressed as the median of the parameter probability distribution, and their uncertainties are estimated using the 16 and 84th percentiles. The resulting SED is shown in Fig. 7. Note that Bagpipes fits the total stellar mass formed. The current stellar mass is , and is also included in Table 3.
We also test more complicated models: first two bursts and then a burst with a delayed -model, assuming Solar metallicities to reduce the number of free parameters. In both cases, the results were very poorly constrained.
In all these fits, we assumed that the central AGN contribution to the flux is minimal. To assess the impact of the AGN emission on the fit, we use the AGN module in Bagpipes (Carnall et al., 2023), which implements a Vanden Berk et al. (2001) quasar model. The model consists of two power laws describing the continuum blueward and redward of 5100 Å in the rest frame, normalised by the flux of the AGN at 5100 Å and by the flux and width of the H emission line. We test different variations of this model, assuming that the AGN is responsible for 10% of the flux at 5100 Å (which we estimate using the most likely SED of the fit without an AGN) and that the flux of the H emission line corresponds to 10% of the band flux. We obtain large reduced values. We then let the 5100 Å normalisation vary and find that the most likely AGN contribution to the flux at 5100 Å is negligible (%) and that the other parameters of the fit correspond to the values listed in Table 3. We thus conclude that the AGN does not contribute significantly to the SED of the BCG complex.
@@
2 The code automatically limits the prior to the age of the Universe at the considered redshift.
@@
4 Bagpipes does not fit directly the current stellar mass, but rather the total stellar mass formed during the star-formation history. We include the current stellar mass for indicative purposes only.
@@ 4.1 Cluster red fraction
The Puddle is a candidate cluster detected in two different surveys, MaDCoWS2 and Euclid (Thongkham et al., 2024a; Euclid Collaboration: Bhargava et al., 2025). It still corresponds to a significant galaxy overdensity after subtraction of the mean galaxy density in the Euclid Deep Fornax Field (see Sect. 3.1). While a single redshift is not enough to spectroscopically confirm a cluster (e.g. Eisenhardt et al., 2008; Noirot et al., 2018), the facts enumerated above suggest that the Puddle cluster is a genuine high-redshift cluster (see also the characterisation of similar unconfirmed clusters by Valtchanov et al. 2004 and Cooke et al. 2016).
As noted in Sect. 3.1, the Puddle cluster red sequence is relatively sparse compared to other clusters and dominated by faint red members. This situation would make the detection of the Puddle cluster challenging with the red-sequence method (e.g. Gladders and Yee, 2000, 2005; Donahue et al., 2002), one of the commonly used algorithms for optical/near-infrared cluster searches (e.g., the C4 clustering algorithm, the Spitzer Adaptation of the Red-sequence Survey and the red-sequence Matched-filter Probabilistic Percolation; Miller et al., 2005; Wilson et al., 2006, 2009; Rykoff et al., 2014).
While Fig. 2 does not show a clear division between the red sequence and the blue cloud, in the colour histograms shown in Fig. 4 bottom right panel, there is a local minimum at . We thus adopt , which corresponds to a stellar age of 1.08 Gyr (Solar metallicity, no attenuation), as our limit between the red sequence and the blue cloud. With this assumption, we find, within a radius centred on the BCG complex, red galaxies () without completeness correction, and red galaxies with it. In both cases, the count of blue galaxies is . These counts correspond to red fractions of % and % respectively. The uncertainties are based on the propagation of the uncertainties of the background subtraction and the completeness correction (see Sect. 3.1.1).
This would make the red fraction of the Puddle cluster consistent with the measurements of Quadri et al. (2012) and Nantais et al. (2016) in clusters, but significantly less quenched than the clusters of Newman et al. (2014), Cooke et al. (2016), and Strazzullo et al. (2019). However, the comparison with other works is made challenging by the different stellar mass, magnitude, and radius cuts applied by each author, and by the dust obscuration, which can be mistaken for quiescence (e.g. Woo et al., 2013; Hatch et al., 2016).
@@ 4.2 BCG formation scenarios
The morphology of the BCG complex is consistent with a complex multiobject merger. While line-of sight superpositions remain possible, the six central cores are embedded in diffuse emission, which suggests that they are interacting. The southwestern core may or may not be part of the merger. The extension of the diffuse emission to the southwest is reminiscent of a tidal tail (see e.g. Webb et al., 2015b). We thus infer that the Puddle cluster is in the process of forming its BCG through a multiobject merger. In this section, we explore the properties of this merger and how it compares with SPT234956 (Miller et al., 2018) and SpARCS104922.6+564032.5 (Webb et al., 2015b). We note two other instances of multiobject mergers interpreted as assembling BCGs in the literature: the Spiderweb protocluster at (Miley et al., 2006) and an unnamed overdensity in the Cosmic Evolution Early Release Science of JWST (Coogan et al., 2023).
@@ 4.2.1 Star formation and AGN activity
Most of the proto-BCGs in the literature exhibit high levels of star formation, obscured or not (Seymour et al., 2012; Webb et al., 2015b; Miller et al., 2018; Coogan et al., 2023). In the Puddle cluster, the SED model that best fits the photometry of the BCG complex is an instantaneous burst of star formation Myr ago, with no more recent star formation. We note however that parametric SED modelling should be taken with caution (e.g. Carnall et al., 2019; Leja et al., 2019; Haskell et al., 2024; Mosleh et al., 2025). Thus, we interpret these results as evidence for a burst of star formation about Myr ago, but do not rule out the possibility of more recent, residual star formation. In fact, given the young age of the dominant stellar population, ongoing unobscured star formation in the Puddle cluster proto-BCG is plausible. Obscured star formation is also possible: all star-formation histories return large dust attenuations, regardless of the metallicity content. Without coverage in the far infrared, it is difficult to assess its importance.
The spectrum of the brightest nucleus does not offer any clues about the current SFR since it is dominated by AGN emission. If we model the spectrum as two Gaussians (see Sect. 3.2), the broadest emission line FWHM is (i.e. ). This is too large to be associated with the merging activity, since the line width is similar to the total velocity dispersion of an entire (proto)cluster (e.g. Kuiper et al., 2011; Ruel et al., 2014).
The results of the three-Gaussian modelling also suggest that the spectrum is AGN-dominated. Using this model, we compute approximate equivalent-width ratios of and . Large values like these are unlikely to be associated with star formation alone, but consistent with the ionisation provided by an AGN (e.g. Bian et al., 2018; Oh et al., 2019, 2022; Garg et al., 2022; Dors et al., 2023; Euclid Collaboration: Scharré et al., 2024; Zhou et al., 2025).
The stellar mass computed by Bagpipes is . This value does not include AGN contamination and assumes a single stellar population, about 300 . An AGN can sometimes result in an overestimation of the stellar mass (e.g. Ciesla et al., 2015; Salim et al., 2016). However, the presence of an older population would instead decrease the mass-to-light ratio (e.g. Mancone and Gonzalez, 2012; Conroy, 2013). Bagpipes stellar mass should thus be treated as an estimate.
@@ 4.2.2 Comparison with SPT234956 simulations
SPT234956 is a protocluster with 14 massive gas-rich galaxies all within kpc of projected distance from each other (Miller et al., 2018). According to the simulations of Rennehan et al. (2020, see also updated versions), SPT234956’s 14 central galaxies should merge into a single BCG in Myr from the time of observation. They find that the star-formation rate peaks very early in their simulation, before decreasing exponentially with a characteristic time of Myr. Remus et al. (2023) comparison with the Magneticum cosmological simulations also predicts a rapid collapse.
@@
Rennehan et al. (2020) also make predictions about the probability of observing massive collapse events like SPT234956 (or the Puddle cluster) at different redshifts. While they observe events at all redshifts, most of them occur before . Yet, observationally, SPT234956 is the highest-redshift merging proto-BCG in the literature (Miller et al., 2018). All the others lie at redshifts closer to the Puddle cluster – between and (Miley et al., 2006; Webb et al., 2015b; Coogan et al., 2023). A sample of five proto-BCGs is not enough to draw statistically significant conclusions. However, the fact that most of them lie at suggests either that most collapses occur later and/or on longer timescales than the Rennehan et al. (2020) predictions, or that selection effects decrease the likelihood of detecting merging BCGs at .
@@
While SPT234956 is the only proto-BCG with dedicated simulations, the SpARCS104922.6+564032.5 system is more directly comparable to the Puddle cluster in terms of redshift and physical scale. At , SpARCS104922.6+564032.5’s star-forming core is dominated by a kpc-long tidal tail, with small clumps of various colours (Webb et al., 2015b). It is unclear if the tidal tail is directly associated with the BCG, or with the pair of smaller merging cluster members located to its north. Subsequent studies have shown that the BCG hosts a faint radio-loud AGN (Trudeau et al., 2019) and that the bulk of the star formation occurs at kpc to the southeast of the BCG (Barfety et al., 2022). This corresponds to the centre of the tidal tail, where the MIPS emission of the system is located. The discovery of X-ray emission centred on the star-forming region suggests that this system might be powered by a displaced cooling flow rather than a merger (Hlavacek-Larrondo et al., 2020).
The presence of a cooling flow in the Puddle cluster cannot be determined without X-ray or SZ observations. However, the Puddle cluster displays none of the clues that suggested the presence of a cooling flow as in SpARCS104922.6+564032.5. Unlike that cluster, the BCG of the Puddle cluster is clearly merging. Furthermore, the centroid of the MIPS emission is almost directly located onto the brightest of the merging galaxies. The proto-BCG emission lines are most likely powered by an AGN (see Sect. 4.2.1) and are inconsistent with star formation as the primary ionising mechanism. The H line is also too broad to be produced by a filamentary nebula similar to those associated with local cooling flows (e.g. McDonald et al., 2010; Hamer et al., 2016; Gaspari et al., 2018; Gendron-Marsolais et al., 2018; Olivares et al., 2019, 2022; Mohapatra et al., 2022). Only the extreme cooling flow in the Phoenix cluster – which is associated with a large amount of unobscured star formation (McDonald et al., 2012, 2013) – produces emission lines with FWHM around (Reefe et al., 2025).

### 13 Mar 2026 v3

@@ 3.3.2 IRAC photometry and morphology modelling
The pixel size of the IRAC mosaics is , which is six times larger than the the pixel size of Euclid VIS and NISP images. To measure the flux of the marginal detections in the IRAC and channels, we resampled our convolved segmentation map to the correct pixel size (setting the pixels with fractional values to unity) and performed the flux and uncertainty measurements as described above.
However, at and the BCG complex is heavily blended with neighbouring galaxies. We used Galfit modelling (Peng et al. 2002, 2010a) to disentangle the structure from neighbouring galaxy fluxes. Using the high resolution of the NISP data (Fazio et al. 2004; Euclid Collaboration: Jahnke et al. 2025), we constructed initial models in band. We first modelled the BCG complex with seven Sérsic profiles for the structure nuclei, and two additional Sérsic profiles for the diffuse emission. Because no PSF estimation is provided with Q1 frames, we approximated Euclid PSFs using a neighbouring unsaturated star. The result is shown in Fig. 6. We then made a separate model of the contaminants before merging the two models together.
Our IRAC and models are directly based on the combined -band models, with a few changes. Unlike the Euclid data, IRAC mosaics have uncertainty maps, which we added into the Galfit settings. We also deleted one of the two diffuse Sérsic profiles, as we could not make the merged models converge otherwise. Likewise, we removed some of the faintest interlopers, as they were too faint to be detected in the IRAC images. For our first pass models, we only let the magnitudes change, but for our final models we allowed positional changes as well to account for shifts induced by PSF miscentring; the effective radii, Sérsic indexes, axis ratios, and position angles were fixed to initial values.
@@
We measured the flux of the BCG complex in the IRAC and channels directly from the best Galfit models. To ensure consistency with the aperture photometry (see Sect. 3.3.1), we computed an aperture correction by dividing the aperture flux of the band by the Galfit flux in the same band. We found an aperture correction of .
@@ 3.3.3 Spectral energy distribution fitting
Table 2 presents the derived DECam, Euclid, and IRAC photometry, which we used to perform spectral energy distribution (SED) modelling of the entire BCG complex. We did not attempt to model the SED of individual cores: The signal-to-noise of the and DECam fluxes was not sufficient to perform a Galfit decomposition.
We adopted most of the Bagpipes (Carnall et al. 2018) standard basic assumptions (models from Bruzual and Charlot 2003, with a Kroupa and Boily 2002 initial mass function, uniform priors), and added a dust component that follows the Calzetti et al. (2000) attenuation law. We replaced the Bagpipes default cosmology with the Planck Collaboration: Aghanim et al. (2020) cosmological parameters. We adopted a parametric approach, testing a delayed -model and a model in which all the stars formed at the same time (i.e. a burst). The burst yields a best fit SED with a smaller reduced than the delayed -model (3.34 compared to 3.75). Furthermore the characteristic time of the delayed -model is extremely short, Myr. The free parameters, priors, and results of the bursty model are presented in Table 3. The results are expressed as the median of the parameter probability distribution, and their uncertainties were estimated using the 16 and 84th percentiles. The resulting SED is shown in Fig. 7. Note that Bagpipes fits the total stellar mass formed. The current stellar mass is , and is also included in Table 3.
We also tested more complicated models: first two bursts and then a burst with a delayed -model, assuming Solar metallicities to reduce the number of free parameters. In both cases, the results were very poorly constrained.
In all these fits, we assumed that the central AGN contribution to the flux is minimal. To assess the impact of the AGN emission on the fit, we used the AGN module in Bagpipes (Carnall et al. 2023), which implements a Vanden Berk et al. (2001) quasar model. The model consists of two power laws describing the continuum blueward and redward of 5100 Å in the rest frame, normalised by the flux of the AGN at 5100 Å and by the flux and width of the H emission line. We tested different variations of this model, assuming that the AGN is responsible for 10% of the flux at 5100 Å (which we estimate using the most likely SED of the fit without an AGN) and that the flux of the H emission line corresponds to 10% of the band flux. We obtained large reduced values. We then let the 5100 Å normalisation vary and found that the most likely AGN contribution to the flux at 5100 Å is negligible (%) and that the other parameters of the fit corresponded to the values listed in Table 3. We thus concluded that the AGN does not contribute significantly to the SED of the BCG complex.
@@
2 The code automatically limited the prior to the age of the Universe at the considered redshift.
@@
4 Bagpipes does not fit directly the current stellar mass, but rather the total stellar mass formed during the star-formation history. We included the current stellar mass for indicative purposes only.
@@ 4.1 Cluster red fraction
The Puddle is a candidate cluster detected in two different surveys, MaDCoWS2 and Euclid (Thongkham et al. 2024a; Euclid Collaboration: Bhargava et al. 2025). It still corresponds to a significant galaxy overdensity after subtraction of the mean galaxy density in the Euclid Deep Fornax Field (see Sect. 3.1). While a single redshift is not enough to spectroscopically confirm a cluster (e.g. Eisenhardt et al. 2008; Noirot et al. 2018), the facts enumerated above suggest that the Puddle cluster is a genuine high-redshift cluster (see also the characterisation of similar unconfirmed clusters by Valtchanov et al. 2004 and Cooke et al. 2016).
As noted in Sect. 3.1, the Puddle cluster red sequence is relatively sparse compared to other clusters and dominated by faint red members. This situation would make the detection of the Puddle cluster challenging with the red-sequence method (e.g. Gladders and Yee 2000, 2005; Donahue et al. 2002), one of the commonly used algorithms for optical/near-infrared cluster searches (e.g. the C4 clustering algorithm, the Spitzer Adaptation of the Red-sequence Survey and the red-sequence Matched-filter Probabilistic Percolation; Miller et al. 2005; Wilson et al. 2006, 2009; Rykoff et al. 2014).
While Fig. 2 does not show a clear division between the red sequence and the blue cloud, in the colour histograms shown in Fig. 4 bottom right panel, there is a local minimum at . We thus adopted , which corresponds to a stellar age of 1.08 Gyr (Solar metallicity, no attenuation), as our limit between the red sequence and the blue cloud. With this assumption, we found, within a radius centred on the BCG complex, red galaxies () without completeness correction, and red galaxies with it. In both cases, the count of blue galaxies was . These counts correspond to red fractions of % and % respectively. The uncertainties were based on the propagation of the uncertainties of the background subtraction and the completeness correction (see Sect. 3.1.1).
This would make the red fraction of the Puddle cluster consistent with the measurements of Quadri et al. (2012) and Nantais et al. (2016) in clusters, but significantly less quenched than the clusters of Newman et al. (2014), Cooke et al. (2016), and Strazzullo et al. (2019). However, the comparison with other works is made challenging by the different stellar mass, magnitude, and radius cuts applied by each author, and by the dust obscuration, which can be mistaken for quiescence (e.g. Woo et al. 2013; Hatch et al. 2016).
@@ 4.2 BCG formation scenarios
The morphology of the BCG complex is consistent with a complex multiobject merger. While line-of sight superpositions remain possible, the six central cores are embedded in diffuse emission, which suggests that they are interacting. The southwestern core may or may not be part of the merger. The extension of the diffuse emission to the southwest is reminiscent of a tidal tail (see e.g. Webb et al. 2015b). We thus inferred that the Puddle cluster is in the process of forming its BCG through a multiobject merger. In this section, we explored the properties of this merger and how it compares with SPT234956 (Miller et al. 2018) and SpARCS104922.6+564032.5 (Webb et al. 2015b). We noted two other instances of multiobject mergers interpreted as assembling BCGs in the literature: the Spiderweb protocluster at (Miley et al. 2006) and an unnamed overdensity in the Cosmic Evolution Early Release Science of JWST (Coogan et al. 2023).
@@ 4.2.1 Star formation and AGN activity
Most of the proto-BCGs in the literature exhibit high levels of star formation, obscured or not (Seymour et al. 2012; Webb et al. 2015b; Miller et al. 2018; Coogan et al. 2023). In the Puddle cluster, the SED model that best fits the photometry of the BCG complex is an instantaneous burst of star formation Myr ago, with no more recent star formation. We note however that parametric SED modelling should be taken with caution (e.g. Carnall et al. 2019; Leja et al. 2019; Haskell et al. 2024; Mosleh et al. 2025). Thus, we interpret these results as evidence for a burst of star formation about Myr ago, but do not rule out the possibility of more recent, residual star formation. In fact, given the young age of the dominant stellar population, ongoing unobscured star formation in the Puddle cluster proto-BCG is plausible. Obscured star formation is also possible: All star-formation histories returned large dust attenuations, regardless of the metallicity content. Without coverage in the far infrared, it is difficult to assess its importance.
The spectrum of the brightest nucleus does not offer any clues about the current SFR since it is dominated by AGN emission. If we model the spectrum as two Gaussians (see Sect. 3.2), the broadest emission line FWHM is (i.e. ). This is too large to be associated with the merging activity, since the line width is similar to the total velocity dispersion of an entire (proto)cluster (e.g. Kuiper et al. 2011; Ruel et al. 2014).
The results of the three-Gaussian modelling also suggest that the spectrum is AGN-dominated. Using this model, we computed approximate equivalent-width ratios of and . Large values like these are unlikely to be associated with star formation alone, but consistent with the ionisation provided by an AGN (e.g. Bian et al. 2018; Oh et al. 2019, 2022; Garg et al. 2022; Dors et al. 2023; Euclid Collaboration: Scharré et al. 2024; Zhou et al. 2025).
The stellar mass computed by Bagpipes is . This value does not include AGN contamination and assumes a single stellar population, about 300 Myr old. An AGN can sometimes result in an overestimation of the stellar mass (e.g. Ciesla et al. 2015; Salim et al. 2016). However, the presence of an older population would instead decrease the mass-to-light ratio (e.g. Mancone and Gonzalez 2012; Conroy 2013). Bagpipes stellar mass should thus be treated as an estimate.
@@ 4.2.2 Comparison with SPT234956 simulations
Lying at , SPT234956 is a protocluster with 14 massive gas-rich galaxies all within kpc of projected distance from each other (Miller et al. 2018). According to the simulations of Rennehan et al. (2020, see also updated versions), SPT234956’s 14 central galaxies should merge into a single BCG in Myr from the time of observation. They found that the star-formation rate peaks very early in their simulation, before decreasing exponentially with a characteristic time of Myr. Remus et al. (2023) comparison with the Magneticum cosmological simulations also predicts a rapid collapse.
@@
Rennehan et al. (2020) also made predictions about the probability of observing massive collapse events like SPT234956 (or the Puddle cluster) at different redshifts. While they observed events at all redshifts, most of them occured before . Yet, observationally, SPT234956 is the highest-redshift merging proto-BCG in the literature (Miller et al. 2018). All the others lie at redshifts closer to the Puddle cluster – between and (Miley et al. 2006; Webb et al. 2015b; Coogan et al. 2023). A sample of five proto-BCGs is not enough to draw statistically significant conclusions. However, the fact that most of them lie at suggests either that most collapses occur later and/or on longer timescales than the Rennehan et al. (2020) predictions, or that selection effects decrease the likelihood of detecting merging BCGs at .
@@
While SPT234956 is the only proto-BCG with dedicated simulations, the SpARCS104922.6+564032.5 system is more directly comparable to the Puddle cluster in terms of redshift and physical scale. At , SpARCS104922.6+564032.5’s star-forming core is dominated by a kpc-long tidal tail, with small clumps of various colours (Webb et al. 2015b). It is unclear if the tidal tail is directly associated with the BCG, or with the pair of smaller merging cluster members located to its north. Subsequent studies have shown that the BCG hosts a faint radio-loud AGN (Trudeau et al. 2019) and that the bulk of the star formation occurs at kpc to the southeast of the BCG (Barfety et al. 2022). This corresponds to the centre of the tidal tail, where the MIPS emission of the system is located. The discovery of X-ray emission centred on the star-forming region suggests that this system might be powered by a displaced cooling flow rather than a merger (Hlavacek-Larrondo et al. 2020).
The presence of a cooling flow in the Puddle cluster cannot be determined without X-ray or SZ observations. However, the Puddle cluster displays none of the clues that suggested the presence of a cooling flow as in SpARCS104922.6+564032.5. Unlike that cluster, the BCG of the Puddle cluster is clearly merging. Furthermore, the centroid of the MIPS emission is almost directly located onto the brightest of the merging galaxies. The proto-BCG emission lines are most likely powered by an AGN (see Sect. 4.2.1) and are inconsistent with star formation as the primary ionising mechanism. The H line is also too broad to be produced by a filamentary nebula similar to those associated with local cooling flows (e.g. McDonald et al. 2010; Hamer et al. 2016; Gaspari et al. 2018; Gendron-Marsolais et al. 2018; Olivares et al. 2019, 2022; Mohapatra et al. 2022). Only the extreme cooling flow in the Phoenix cluster – which is associated with a large amount of unobscured star formation (McDonald et al. 2012, 2013) – produces emission lines with FWHM around (Reefe et al. 2025).